rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#2813 — FS#6773 — rbx-97-n7

Attached to Project— Dedicated servers
Modernization
all (dedicated servers)
CLOSED
100%
We found a problem of routing table on a card of rbx-97-n7.

2012 May 18 16:38:20 rbx-97-n7 %IPFIB-SLOT3-2-FIB_TCAM_RESOURCE_EXHAUSTION: FIB TCAM exhausted for IPV4 unicast
2012 May 18 16:38:20 rbx-97-n7 %IPFIB-SLOT3-4-FIB_TCAM_AF_INSERT_FAIL: FIB TCAM prefix insertion fail for IPV4 unicast
2012 May 18 16:38:36 rbx-97-n7 %IPFIB-SLOT3-2-FIB_TCAM_RESOURCE_EXHAUSTION: FIB TCAM exhausted for IPV4 unicast
2012 May 18 16:38:36 rbx-97-n7 %IPFIB-SLOT3-4-FIB_TCAM_AF_INSERT_FAIL: FIB TCAM prefix insertion fail for IPV4 unicast
2012 May 18 16:38:53 rbx-97-n7 %IPFIB-SLOT3-2-FIB_TCAM_RESOURCE_EXHAUSTION: FIB TCAM exhausted for IPV4 unicast
2012 May 18 16:38:53 rbx-97-n7 %IPFIB-SLOT3-4-FIB_TCAM_AF_INSERT_FAIL: FIB TCAM prefix insertion fail for IPV4 unicast
rbx-97-n7#

In fact this card is in configuration "light" and does not possess enough resources for routing.

rbx-97-n7# sh mod
Mod Ports Module-Type Model Status
--- ----- ----------------------------------- ------------------ ----------

3 32 10 Gbps Ethernet Module N7K-M132XP-12 ok

We should switch to a new card. A cut-off of 5 to 10 seconds on the NAS service is expected.

We will begin the maintenance on the night of Wednesday, May 23rd to Thursday, May 24th at 0:00 (EST).
Date:  Friday, 25 May 2012, 02:47AM
Reason for closing:  Done
Comment by OVH - Thursday, 24 May 2012, 01:37AM

We started.


Comment by OVH - Thursday, 24 May 2012, 01:39AM

A pair of N5 has been switched. We have a problem connecting to the 2nd pair. We are investigating.


Comment by OVH - Thursday, 24 May 2012, 01:40AM

A vPC problem on the N5 was the reason. After having corrected the configuration, the ports are up.

There was a cut-off of about 5 min.